OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2.

hartnn (hartnn):

the point slope form of equation of line is \(\large y-y_1=m(x-x_1)\) where m=slope =-2 here. point =(x1,y1)= (3,0) --->x1=3,y1=0 just plug in values!

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

I put Y=-2X as the answer and it said I was wrong :(

hartnn (hartnn):

because its incorrect...didn't you get what i have written ?? you just need to plug in values in that equation.

OpenStudy (anonymous):

I'm not good at math lol

hartnn (hartnn):